xy + 2xy = 6 of lvpolis yozga Senes Youtnjos -gla)=3 o Power
Answer #1

Q. Here Let the solution of the differential equation ( 12+1) y + ny + 2ny = 0 0 19 y 2 N=0 annn an. nnn-1 n=0 So, dn d24 a► 202 + a canna + ant2 (n+2)(n+1) +2 an-1) 22 n=1 comparing both side we get Aq=0. anna + ant2 (n+2)(n +1) +20n-1=0 + On+2 -So, a6 = + 8 ai 2aotai 3 6.5 200 +901 3.6.5 D A G - 20o-al 3.2 + 35 + 3.6.5 ai ain3 6 40 + ao 415 10 40 :-) + 26 + 45 + 3 n 5Check the last box... this is your series solution...